OpenStudy (automaticloveletter):

A system of equations is shown below: n = 3m + 5 n − 2m = 3 What is the solution, in the form (m, n), to the system of equations? (2, 11) (1, 8) (−2, −1) (−3, −4)

Mehek (mehek14):

what do you think?

OpenStudy (automaticloveletter):

No clue

Mehek (mehek14):

let's substitute the first equation into the second one \(3m + 5 -2m = 3 \\ 5 - m = 3 \\ -m = -2 \\ m = 2\)

Mehek (mehek14):

substitute 2 for m \(n=3*2+5\\3*2=6\\6+5=11\)

Mehek (mehek14):

so (2,11)

OpenStudy (automaticloveletter):

The system of equations shown below is graphed on a coordinate grid: 3y + x = 4 2y − x = 6 Which statement is true about the coordinates of the point that is the solution to the system of equations? It is (−2, 2) and lies on both lines. It is (−5, 3) and lies on both lines. It is (−5, 3) and does not lie on either line. It is (−2, 2) and does not lie on either line.

Mehek (mehek14):

actually it's -2,-1 when I graphed it

Mehek (mehek14):

and for this one it's A